OpenStudy (anonymous):

2(W^2 + 14W - 378) = 0

OpenStudy (anonymous):

Can you use the Delta formula ? \[\Delta=b^2-4ac~~~x=\frac{-b\pm\sqrt\Delta}{2a}\] or not ?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

The rectangular dance floor is twice as long as it is wide. The tent surrounds the dance floor, leaving some space (7 feet in each direction) for guests to mingle and cool off after dancing. This extra space (represented by the grey area) has an area of 952 square feet. Your task is to find the dimensions of the dance floor. Be sure to show all work. Dimensions Outside:Top= 2x+14 Side=x+14 Gray Area= 952 ft.^2 Dimensions Dance Floor:Top=2x Side= x

OpenStudy (anonymous):

I need to find the dimensions of the dance floor.

OpenStudy (anonymous):

u have a pic of this right? could u post it?

OpenStudy (anonymous):

the area total is represented by (x+14)(2x+14) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and the area could bbe represented by 2x(x) + 952 too is that right for u?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

so u need to make equal both expresions and solve for x

OpenStudy (anonymous):

Do you that that the quadratic formula is \[x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\] where a, b, and c are from ax^2+bx+c=0